PS1 IPP Czar Logs for the week YYYY.MM.DD - YYYY.MM.DD

(Up to PS1 IPP Czar Logs)

Monday : 2017.06.05

  • MEH: again fixing stalled stamps for QUB

Tuesday : 2017.06.06

  • MEH: Heather wondering issue w/ ippx012 being down -- ganglia shows offline early yesterday, console shows
    <Jun/05 06:14 am>ippx012 login: [1679460.508553] ------------[ cut here ]------------
    <Jun/05 06:14 am>[1679470.992940] kernel BUG at kernel/timer.c:926!
    <Jun/05 06:14 am>[1679471.414703] invalid opcode: 0000 [#1] SMP 
    <Jun/05 06:15 am>[1679473.103464] Modules linked in: adm1021 ipv6 dm_mod joydev usbhid microcode igb ehci_hcd uhci_hcd i2c_i801 ioatdma coretemp i2c_core pcspkr usbcore acpi_cpufreq mperf dca usb_common sg processor button thermal_sys
    <Jun/05 06:15 am>[1679506.734049] CPU 10 
    <Jun/05 06:15 am>[1679506.736226] Pid: 14368, comm: apache2 Not tainted 3.7.6 #1 Supermicro X8DTT/X8DTT
    <Jun/05 06:15 am>[1679508.835724] RIP: 0010:[<ffffffff8103882d>]  [<ffffffff8103882d>] add_timer_on+0x34/0x78
    <Jun/05 06:15 am>[1679510.520854] RSP: 0018:ffff88183f343e10  EFLAGS: 00010286
    
    • no remote power management yet, Haydn reminds would reboot 4 nodes ippx009-x012 and Heather would prefer to wait until next week to just individually reboot w/ physical button press
  • MEH: more fixing of error_cleaned data so pstamp and updates can work
  • MEH: ippc18/homedir was quickly running out of space -- Heather cleaned up logs to make space for ongoing, regular batch running, the warning cron (ipp@ippc18) has been enabled to alert when gets to minimal levels
  • MEH: K2.nightlyscience processing chip-cam-warp automated again and should be checked on by czars like the rest of nightly processing for next month or so again -- WSdiffs are still managed externally by me

Wednesday : 2017.06.07

  • MEH: another case of nightly_science.pl not doing alternate visit diffims for MOPS -- v2-v3 manually queued
    | exp_id  | exp_name    | filter  | fault | fwhm_major | quality | dateobs             | comment                               |
    +---------+-------------+---------+-------+------------+---------+---------------------+---------------------------------------+
    | 1255359 | o7911g0410o | i.00000 |     0 |    5.33247 |       0 | 2017-06-07 12:15:44 | OSSR.R21N2.17.Q.i ps1_33_0247 visit 1 | 
    | 1255377 | o7911g0428o | i.00000 |     0 |    5.20372 |       0 | 2017-06-07 12:32:56 | OSSR.R21N2.17.Q.i ps1_33_0247 visit 2 | 
    | 1255395 | o7911g0446o | i.00000 |     0 |    5.44651 |       0 | 2017-06-07 12:50:08 | OSSR.R21N2.17.Q.i ps1_33_0247 visit 3 | 
    | 1255413 | o7911g0464o | i.00000 |     0 |    5.09305 |    4007 | 2017-06-07 13:07:15 | OSSR.R21N2.17.Q.i ps1_33_0247 visit 4 | 
    

Thursday : YYYY.MM.DD

Friday : 2017.06.09

  • MEH: MOPS test diffs running on ipps nodes
  • 21:00 EAM : restarted the pantasks

Saturday : YYYY.MM.DD

Sunday : YYYY.MM.DD